OpenStudy (anonymous):

which equation represents the line that contains the point (-2,9) and is perpendicular to the line y=2x+6? A.y=-1/2x-7 B.y=-1/2x+8 C.y=-1/2x+11 D. y=2x+13

OpenStudy (anonymous):

help please

OpenStudy (mww):

You can use the point-gradient formula. \[y - y1 = m(x -x1)\] The point would be the one you were given and the gradient is the one that would make it perpendicular to y = 2x + 6. (Use the fact that \[m1 \times m2 = -1\] if the lines are perpendicular.) Once you find the gradient, plug in all the vaues into the formula.

OpenStudy (anonymous):

your looking for a perpendicular line which means the slope will be the negative inverse of the slope of your given equation. Since that slope is 2, then the perpendicular slope will be -1/2. Now we use y - y1 = m(x - x1) using slope -1/2 and points (-2,9) lets sub our info in... y - 9 = -1/2(x -(-2) y - 9 = -1/2x + 2 y = -1/2x + 2 + 9 y = -1/2x + 11 <-- answer C